# An Information Bottleneck Approach for Markov Model Construction

### v1

@@ I Introduction
A powerful and popular analysis method to study the dynamical behavior of any molecular system displaying a sufficiently level of complexity is through Markov state models (MSMs)[5, 6, 7]. MSMs can not only predict long-timescale dynamics using multiple short MD trajectories but also serve as a bridge between high-resolution MD simulations and a more macroscopic description of the dynamics.[3, 8, 9] This makes them a powerful tool for understanding kinetic processes in complex molecular systems.
@@
A number of approaches leveraging artificial neural networks have demonstrated efficacy within the context of variational optimization workflows. Recently, VAMPnets, which employ the VAMP and Koopman theory as their guiding principle, have been developed as a single end-to-end data-driven model to directly map molecular coordinates to coarse-grained macrostates.[10] Later, a state-free reversible VAMPnets (SRVs) is specifically introduced to learn nonlinear approximates to the leading slow eigenfunctions[18] and has also been generalized to biased simulations[19, 20]. Time-lagged autoencoders utilize auto-associative neural networks to reconstruct a time-lagged signal, and embeded variables extracted from the bottleneck layer can serve as the CVs.[21] Another approach, variational dynamics encoders (VDEs), combine time-lagged reconstruction loss and autocorrelation maximization within a variational autoencoder to approximate the dynamical propagator.[22]
In this work, we demonstrate a robust protocol for MSMs construction through the use of an information bottleneck approach called State Predictive Information Bottleneck (SPIB).[23] In previous studies[24, 25, 26, 27, 28, 29, 30], SPIB learns useful low dimensional CVs for enhanced sampling in molecular simulations to speed up the diverse processes ranging from permeation and dissociation of medically relevant ligands, conformational changes in proteins to nucleation of crystal polymorphs. Here we focus on demonstrating SPIB as a state-of-the-art approach for automatic construction of multi-resolution MSMs. Different from the existing methods, such as VAMP-based methodologies, which maximizes the Rayleigh coefficient or VAMP score, SPIB integrates the information bottleneck framework with a simple heuristic of the state metastability at a pre-specified lag time to achieve feature extraction and state division in a unified approach. By leveraging this lag time parameter, SPIB facilitates adaptive adjustment of the number of metastable states within the final kinetic model, enabling the automatic generation of MSMs at different resolutions. As a result, SPIB facilitates the automatic clustering and projection of the MD simulation data into a few macrostates and learns directly from MD trajectories a low dimensional latent space where these states are cleanly separated into metastable states, providing a single end-to-end framework that integrates dimension reduction, clustering, and lumping tasks.
@@ II.1 Markov State Model (MSM)
Let $\mathbf{X}\subseteq\mathbb{R}^{3N}$ denote a high-dimensional signal characterizing the configuration of a generic molecular system. Then a MSM is constructed by clustering $\mathbf{X}$ into a set of states $\mathbf{y}=\left\{1,2,3,\ldots,S\right\}$. For the folding of mini-proteins, $\mathbf{X}$ could be expressed in terms of internal coordinates, such as the minimal residue-residue distances. The states $\mathbf{y}$ could comprise the folded state, unfolded state, misfolded state and some important intermediate states. Let the values measured at time $t$ be denoted by $\mathbf{X}_{t}$ and $\mathbf{y}_{t}$ respectively. In this way, the trajectory $\{\mathbf{X}_{t}\}_{t=0}^{T}$ is coarse-grained to a discrete trajectory $\{\mathbf{y}_{t}\}_{t=0}^{T}$ that indicates which state the system visits at time t. Using statistics from $\mathbf{y}_{t}$, we can calculate the conditional probabilities of moving between each of the discrete states between a user-defined lag time $\tau$ and store the conditional probabilities in a transition probability matrix (TPM) $\mathbf{T}(\tau)$. If the model is validated as Markovian, the long-timescale dynamics could be modeled through the first-order master equation: $\mathbf{T}(n\tau)=\mathbf{T}^{n}(\tau)$. Both thermodynamic properties (e.g. stationary state populations) and kinetic (e.g mean first passage time (MFPT), transition pathways and implied timescales (ITS)) can then be calculated from the TPM. Specifically, if the TPM is diagonalized with the left and right eigenvectors and eigenvalues: $T_{ij}(\tau)=\phi^{T}_{i}(\tau)\lambda_{i}(\tau)\psi_{j}(\tau)$, the stationary populations can be directly obtained from the left eigenvector $\phi_{1}(\tau)$ (if the TPM is row-normalized) with eigenvalues equal to 1. The remaining eigenvectors can be utilized to infer the transition mechanisms between the states. And the ITS for different eigenmodes can be defined as $t_{i}(\tau)=\frac{-\tau}{\ln(\lambda_{i}(\tau))}$. We refer to Refs. 3, 8, and 13 for an excellent recent technical review of MSMs.
@@ II.2 Variational approach for Markov processes (VAMP)
The variational approach for Markov processes (VAMP), developed by Wu and Noé [48, 49], provides a variational framework to approximate the dynamical propagator of general Markovian dynamical processes, whether reversible or irreversible, stationary or non-stationary. The VAMP theory introduces the VAMP-r score, defined as the sum of the singular values raised to the power $r$ of the approximated propagator, to quantify the difference between the approximated leading singular functions of the dynamical propagator and the ground truth. A higher score indicates a superior approximation of the singular functions which represent the leading slowest dynamical modes of the system. Based on the this, VAMPnets have been developed[10], emerging as an end-to-end, neural-network-based, unsupervised method for constructing MSMs. VAMPnets can directly map internal coordinate features to macrostate assignments probabilities, thereby replacing the intricate intermediate pipeline. Additionally, if the studied dynamics are time-reversible and detailed balanced, a specialized variational approach for conformational dynamics (VAC) can be derived.[50, 17]
@@ II.3 State Predictive Information Bottleneck (SPIB)
where $K$ is the number of representative-inputs, and $\omega_{k}$ represents the weight of $p_{\theta}(\mathbf{z}|\mathbf{X}_{\text{rep}}^{k})$ under the constraint $\sum_{k}\omega_{k}=1$. The algorithm to determine the representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$ operates as follows: Initially, one sample is randomly selected for each initial state to form our initial set of representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$, where $K$ corresponds to the number of initial states. After each iteration of model training and state label refinement, all input samples are mapped to the learned latent space. We compute the center of each newly refined, non-empty metastable state in the learned latent space and identify the nearest sample for each center using Euclidean distance in the latent space. These selected samples then serve as the new representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$ in the subsequent iteration. Consequently, our algorithm automatically adjusts the representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$, and the number of representative inputs $K$ always matches the number of states in $\mathbf{y}$. Moreover, by incorporating the mixture of Gaussians prior in Eq. LABEL:eq:SPIB_obj, the regularization term is able to encourage spatial separation among the state centers within the latent space.
@@ II.4 Baseline models and Quantitative Metrics
    $\text{GMRQ}=\sum_{i=1}^{n}\lambda_{i},$
@@
    Metastability: defined as the mean of the trace of the TPM, which measures how probable it is for the system to remain in the same state after a lag time $\tau$ parametrizing the TPM. Metastability is a useful metric for evaluating the model, and high metastability typically indicates effective separation of slow inter-state dynamics from fast intra-state dynamics.[7]
@@
    $\text{DBI}=\frac{1}{N}\sum_{i}\max_{j}\frac{s_{i}+s_{j}}{r_{ij}},$
@@ II.5 Systems, MSM Construction and Validation
All three datasets are featurized with all minimal residue-residue distances (calculated as the closest distance between the heavy atoms of two residues separated in sequence by at least two neighboring residues), resulting in 153 features for Trp-cage, 528 features for both HP35 and WW-domain. For computational feasibility, we analyze the HP35 and WW-domain at a resolution 1 ns per frame, but for the smaller Trp-cage system we retain the 0.2 ns per frame resolution of the original trajectory.
@@
To prepare the data, we first divide the long equilibrium DESRES trajectory into 100 equal-length short segments. These segments are treated as independent trajectories, mimicking the practical scenario of MSM construction where multiple short trajectories are collected. For cross-validation, the segments are then shuffled and subsampled as part of the train-test split procedure for each fold. All parameters are then scored by considering only the top 3 eigenvalues at a 100 ns MSM lag time for all three systems. The 100ns lag time is validated as Markovian lag time for all three systems based on the ITS analysis and CK test (see SI). All GRMQ scores and metastability values reported in this paper are calculated using MSMs with 100ns lag time.
@@ III.2 Quantitative assessment
To achieve a more coarse-grained representation of the SPIB states, we tune $\Delta t$ value to yield, on average, between 4 to 6 metastable states. Specifically, for Trp-cage, we select $\Delta t=100$ ns which yielded 4 metastable states using SPIB. For HP35, we use $\Delta t=110$ ns, which generates models with an average of 5 metastable states. Finally, for the WW-domain, we use $\Delta t=70$ ns which led to 5 metastable states with SPIB. We note that, given the results presented in Figure 3, different lag times within a small range of the selected values for each protein will yield qualitatively similar SPIB models. In this section, we show how these coarse SPIB models can be used as kinetic models competitive or superior to other state-of-the-art methods for MSM construction. These high-performance kinetic models constructed using SPIB are generated without explicitly optimizing along the measured performance metrics and are robust across a wide range of hyperparameter space, as shown in gruesome detail in the SI.
@@
Across all three systems, SPIB outperforms tICA-PCCA+ and PCA-PCCA+ methods in the ITS, as the latter two fail to converge at the same lag times, and the converged timescales are slightly smaller. This observed trend is attributed to the capacity of SPIB to learn nonlinear transformations of input coordinates, providing enhanced resolution of slower processes, and the use of a continuous basis set for MSM macrostates construction. These enhanced capabilities yeild a reduction of discretization errors compared to the Galerkin method when approximating the dynamical propagator. This results in higher converged timescales and the generation of state models with clearer time separations, leading to shorter relaxation times within states and, consequently, a reduced Markovian lag time.
Table 1: Quantitative comparison of different methods for large $\Delta t$ in terms of GMRQ (scoring based on the top 3 eigenvalues), metastability, Shannon entropy, and DBI across three systems. The reported values represent the mean along with the standard error of the mean derived from 10-fold cross-validation results.
@@
Table 2: Quantitative comparison of different methods for moderate $\Delta t$ in terms of GMRQ (scoring based on the top 5 eigenvalues), metastability, Shannon entropy, and DBI across three systems. The reported values represent the mean along with the standard error of the mean derived from 10-fold cross-validation results.
@@ III.3.1 Trp-Cage
Figure 5: Qualitative description of the MSM analysis for Trp-cage protein. (a) and (b) give the free-energy surfaces in the two-dimensional SPIB latent space for large and moderate $\Delta t$, respectively. (c) and (d) give the metastable states learned by SPIB in the case of large and moderate $\Delta t$, respectively. (e) The Sankey plot illustrates the corresponding relations between states learned by SPIB using large (left) and moderate (right) $\Delta t$. (f) The MSM constructed based on states identified by SPIB, trained with a moderate $\Delta t$, is visualized using a flux network. The node size is proportional to the stationary population of the states, and the arrow width is scaled according to jump probabilities. Additionally, ten randomly selected conformations from each state are overlapped and displayed adjacent to the corresponding node.
@@ III.3.2 HP35
The other 8 states in the 10 state model correspond to states with various degrees of unfolding or misfolded states. Unfolded state 1 in the 5 state model is decomposed into states S1, S2, S3, S4, and S7 in the 10 state model. All these unfolded states contain some degree of secondary structure, but differ the fraction and location of the structure. Specifically, state S1 presents with alpha helix 1 ($\alpha_{1}$) folded, $\alpha_{2}$ unfolded, and $\alpha_{3}$ misfolded, S2 with $\alpha_{1}$ unfolded and the other two helices misfolded, S3 and S4 with $\alpha_{1}$ unfolded and the other two helices folded, and S7 presenting as a completely unfolded state.
@@
Generally, S7 appears to be a sink for both the misfolded state S1 and the other unfolded states. However, overall, the qualitative analysis of the 10 state model indicates that there is not a dominant folding-unfolding pathway for the HP35 protein; it instead indicates that, even for this simple protein, the folding processes is multifaceted without a single dominant route.
A comparison of the SPIB metastable states at both the 5 and 10 state level of resolution and those discovered by the PCCA+, MPP, and VAMPnet approaches are given in the SI. There we observe that only SPIB and tICA-MPP are able to distinguish clearly the two folded states discriminated by the $\phi_{3}$ dihedral while all other methodologies tested lump those two states into one metastable, folded state at the coarser 5 state level. When the clustering is performed at the finer 10 state level, all methods are able to distinguish the two folded states differing in $\phi_{3}$ angle, although the VAMPnet approach only resolves the two folded states for 40% of the models built. This result is consistent with VAMPnet’s occasional failure in predicting the third-slowest timescale in an alanine dipeptide system[10].

### v2

@@ I Introduction
A powerful and popular analysis method to study the dynamical behavior of any molecular system displaying a sufficient level of complexity is through Markov state models (MSMs)[5, 6, 7]. MSMs can not only predict long-timescale dynamics using multiple short MD trajectories but also serve as a bridge between high-resolution MD simulations and a more macroscopic description of the dynamics.[3, 8, 9] This makes them a powerful tool for understanding kinetic processes in complex molecular systems.
@@
A number of approaches leveraging artificial neural networks have demonstrated efficacy within the context of variational optimization workflows. Recently, VAMPnets, which employ the VAMP and Koopman theory as their guiding principle, have been developed as a single end-to-end data-driven model to directly map molecular coordinates to coarse-grained macrostates.[10] Later, a state-free reversible VAMPnets (SRVs) is specifically introduced to learn nonlinear approximates to the leading slow eigenfunctions[18] and has also been generalized to biased simulations[19, 20]. Time-lagged autoencoders utilize auto-associative neural networks to reconstruct a time-lagged signal, and embedded variables extracted from the bottleneck layer can serve as the CVs.[21] Another approach, variational dynamics encoders (VDEs), combine time-lagged reconstruction loss and autocorrelation maximization within a variational autoencoder to approximate the dynamical propagator.[22]
In this work, we demonstrate a robust protocol for MSMs construction through the use of an information bottleneck approach called State Predictive Information Bottleneck (SPIB).[23] In previous studies, SPIB learns useful low dimensional CVs for enhanced sampling in molecular simulations to speed up the diverse processes ranging from permeation[24] and dissociation[30, lee2024calculating] of medically relevant ligands, conformational changes in proteins[28, 29] to nucleation of crystal polymorphs[26, 27]. Here, we focus on demonstrating SPIB as a state-of-the-art approach for automatic construction of multi-resolution MSMs. Different from the existing methods, such as VAMP-based methodologies, which maximizes the Rayleigh coefficient or VAMP score, SPIB integrates the information bottleneck framework with a simple heuristic of the state metastability at a pre-specified lag time to achieve feature extraction and state division in a unified approach. By leveraging this lag time parameter, SPIB facilitates adaptive adjustment of the number of metastable states within the final kinetic model, enabling the automatic generation of MSMs at different resolutions. As a result, SPIB facilitates the automatic clustering and projection of the MD simulation data into a few macrostates and learns directly from MD trajectories a low dimensional latent space where these states are cleanly separated into metastable states, providing a single end-to-end framework that integrates dimension reduction, clustering, and lumping tasks.
@@ II.1 Markov State Model (MSM)
Let $\mathbf{X}\subseteq\mathbb{R}^{3N}$ denote a high-dimensional signal characterizing the configuration of a generic molecular system. Then a MSM is constructed by clustering $\mathbf{X}$ into a set of states $\mathbf{y}=\left\{1,2,3,\ldots,S\right\}$. For the folding of mini-proteins, $\mathbf{X}$ could be expressed in terms of internal coordinates, such as the minimal residue-residue distances. The states $\mathbf{y}$ could comprise the folded state, unfolded state, misfolded state and some important intermediate states. Let the values measured at time $t$ be denoted by $\mathbf{X}_{t}$ and $\mathbf{y}_{t}$ respectively. In this way, the trajectory $\{\mathbf{X}_{t}\}_{t=0}^{T}$ is coarse-grained to a discrete trajectory $\{\mathbf{y}_{t}\}_{t=0}^{T}$ that indicates which state the system visits at time t. Using statistics from $\mathbf{y}_{t}$, we can calculate the conditional probabilities of moving between each of the discrete states between a user-defined lag time $\tau$ and store the conditional probabilities in a transition probability matrix (TPM) $\mathbf{T}(\tau)$. If the model is validated as Markovian, the long-timescale dynamics could be modeled through the first-order master equation: $\mathbf{T}(n\tau)=\mathbf{T}^{n}(\tau)$. Both thermodynamic properties (e.g. stationary state populations) and kinetic (e.g mean first passage time (MFPT), transition pathways and implied timescales (ITS)) can then be calculated from the TPM. Specifically, if the TPM is diagonalized with the left and right eigenvectors and eigenvalues: $T_{ij}(\tau)=\phi^{T}_{i}(\tau)\lambda_{i}(\tau)\psi_{j}(\tau)$, the stationary populations can be directly obtained from the left eigenvector $\phi_{1}(\tau)$ (if the TPM is row-normalized) with eigenvalues equal to 1. The remaining eigenvectors can be utilized to infer the transition mechanisms between the states. And the ITS for different eigenmodes can be defined as $t_{i}(\tau)=\frac{-\tau}{\ln(\lambda_{i}(\tau))}$. We refer to Refs. 3, 8, 13 for an excellent recent technical review of MSMs.
@@ II.2 Variational approach for Markov processes (VAMP)
The variational approach for Markov processes (VAMP), developed by Wu and Noé [48, 49], provides a variational framework to approximate the dynamical propagator of general Markovian dynamical processes, whether reversible or irreversible, stationary or non-stationary. The VAMP theory introduces the VAMP-r score, defined as the sum of the singular values raised to the power $r$ of the approximated propagator, to quantify the difference between the approximated leading singular functions of the dynamical propagator and the ground truth. A higher score indicates a superior approximation of the singular functions which represent the leading slowest dynamical modes of the system. Based on this, VAMPnets have been developed[10], emerging as an end-to-end, neural-network-based, unsupervised method for constructing MSMs. VAMPnets can directly map internal coordinate features to macrostate assignments probabilities, thereby replacing the intricate intermediate pipeline. Additionally, if the studied dynamics are time-reversible and detailed balanced, a specialized variational approach for conformational dynamics (VAC) can be derived.[50, 17]
@@ II.3 State Predictive Information Bottleneck (SPIB)
where $K$ is the number of representative-inputs, and $\omega_{k}$ represents the weight of $p_{\theta}(\mathbf{z}|\mathbf{X}_{\text{rep}}^{k})$ under the constraint $\sum_{k}\omega_{k}=1$. The algorithm to determine the representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$ operates as follows: Initially, one sample is randomly selected for each initial state to form our initial set of representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$, where $K$ corresponds to the number of initial states. After each iteration of model training and state label refinement, all input samples are mapped to the learned latent space. We compute the center of each newly refined, non-empty metastable state in the learned latent space and identify the nearest sample for each center using Euclidean distance in the latent space. These selected samples then serve as the new representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$ in the subsequent iteration. Consequently, our algorithm automatically adjusts the representative inputs $\{\mathbf{X}_{\text{rep}}^{k}\}_{k=1}^{K}$, and the number of representative inputs $K$ always matches the number of states in $\mathbf{y}$. Moreover, by incorporating the mixture of Gaussians prior to Eq. LABEL:eq:SPIB_obj, the regularization term is able to encourage spatial separation among the state centers within the latent space.
@@ II.4 Baseline models and Quantitative Metrics
    $\text{GMRQ}=\sum_{i=1}^{n}\lambda_{i},$ (8)
@@
    Metastability $Q$: defined as the mean of the trace of the TPM, which measures how probable it is for the system to remain in the same state after a lag time $\tau$ parametrizing the TPM. Metastability is a useful metric for evaluating the model, and high metastability typically indicates effective separation of slow inter-state dynamics from fast intra-state dynamics.[7] Explicitly, we define the metastability $Q$ as
    $Q=\frac{1}{S}\text{tr}\left(\mathbf{T}(\tau)\right),$ (9)
    where $S$ is, as previously defined, the number of metastable states spanned by $\mathbf{T}(\tau)$.
@@
    $\text{DBI}=\frac{1}{N}\sum_{i}\max_{j}\frac{s_{i}+s_{j}}{r_{ij}},$ (11)
@@ II.5 Systems, MSM Construction and Validation
All three datasets are featurized with all minimal residue-residue distances (calculated as the closest distance between the heavy atoms of two residues separated in sequence by at least two neighboring residues), resulting in 153 features for Trp-cage, 528 features for both HP35 and WW-domain. The features constructed from distances between the closest heavy atoms offer the advantage of accurately depicting the formation and breaking of local residue contacts, thereby facilitating the identification of structurally well-defined conformational states. But we note that these embedded features may result in non-differentiable CVs due to the switching of closest atom pairs. Therefore, utilizing distances between $C_{\alpha}$ atoms or mass centers may yield CVs more appropriate for enhanced sampling. For computational feasibility, we analyze the HP35 and WW-domain at a resolution 1 ns per frame, but for the smaller Trp-cage system we retain the 0.2 ns per frame resolution of the original trajectory.
@@
To prepare the data, we first divide the long equilibrium DESRES trajectory into 100 equal-length short segments. These segments are treated as independent trajectories, mimicking the practical scenario of MSM construction where multiple short trajectories are collected. For 10-fold cross-validation, the 100 segments are then shuffled and randomly subsampled as part of the train-test split procedure for each fold. This straightforward data processing method has been employed in the previous study[60]. Given the sufficiently long protein folding trajectories provided by DESRES, we anticipate that this simple train-test split procedure will adequately ensure that each split represents the dynamics of interest. All parameters are then scored by considering only the top 3 eigenvalues at a 100 ns MSM lag time for all three systems. The 100ns lag time is validated as Markovian lag time for all three systems based on the ITS analysis and CK test (see SI). All GRMQ scores and metastability values reported in this paper are calculated using MSMs with 100 ns lag time.
@@ III.2 Quantitative assessment
To achieve a more coarse-grained representation of the SPIB states, we tune $\Delta t$ value to yield, on average, 4 or 5 metastable states. Specifically, for Trp-cage, we select $\Delta t=100$ ns which yielded 4 metastable states using SPIB. For HP35, we use $\Delta t=110$ ns, which generates models with an average of 5 metastable states. Finally, for the WW-domain, we use $\Delta t=70$ ns which led to 5 metastable states with SPIB. We note that, given the results presented in Figure 3, different lag times within a small range of the selected values for each protein will yield qualitatively similar SPIB models. In this section, we show how these coarse SPIB models can be used as kinetic models competitive or superior to other state-of-the-art methods for MSM construction. These high-performance kinetic models constructed using SPIB are generated without explicitly optimizing along the measured performance metrics and are robust across a wide range of hyperparameter space, as shown in gruesome detail in the SI.
@@
For the SPIB models examined, the dimensionality of the SPIB latent space used to construct MSMs is always two-dimensional. For the PCA and tICA approaches, the GMRQ is maximized when using a higher dimensionality. For PCA, we utilized the 10, 10, and 20 PCs with the highest variances for constructing MSMs for Trp-cage, HP35, and WW-domain, respectively. For Trp-cage, HP35, and WW-domain, we use 3, 4, and 6 tICA components for MSM construction, respectively. The GMRQ scores for MSMs constructed on PCA, tICA, and SPIB spaces of varying dimensionality are given in the SI.
@@
Across all three systems, SPIB outperforms tICA-PCCA+ and PCA-PCCA+ methods in the ITS, as the latter two fail to converge at the same lag times, and the converged timescales are slightly smaller. This observed trend is attributed to the capacity of SPIB to learn nonlinear transformations of input coordinates, providing enhanced resolution of slower processes, and the use of a continuous basis set for MSM macrostates construction. These enhanced capabilities yield a reduction of discretization errors compared to the Galerkin method when approximating the dynamical propagator. This results in higher converged timescales and the generation of state models with clearer time separations, leading to shorter relaxation times within states and, consequently, a reduced Markovian lag time.
Table 1: Quantitative comparison of different methods for large $\Delta t$ in terms of GMRQ (scoring based on the top 3 eigenvalues), metastability $Q$, Shannon entropy, and DBI across three systems. The arrows indicate whether larger or smaller values are better for each metric. The reported values represent the mean along with the standard error of the mean derived from 10-fold cross-validation results.
@@
Table 2: Quantitative comparison of different methods for moderate $\Delta t$ in terms of GMRQ (scoring based on the top 5 eigenvalues), metastability $Q$, Shannon entropy, and DBI across three systems. The reported values represent the mean along with the standard error of the mean derived from 10-fold cross-validation results.
@@ III.3.1 Trp-Cage
Figure 5: Qualitative description of the MSM analysis for Trp-cage protein. (a) and (b) give the free-energy surfaces in the two-dimensional SPIB latent space, denoted by $\text{IB}_{0}$ and $\text{IB}_{1}$, for large and moderate $\Delta t$, respectively. (c) and (d) give the metastable states learned by SPIB in the case of large and moderate $\Delta t$, respectively. (e) The Sankey plot illustrates the corresponding relations between states learned by SPIB using large (left) and moderate (right) $\Delta t$. (f) The MSM constructed based on states identified by SPIB, trained with a moderate $\Delta t$, is visualized using a flux network. The node size is proportional to the stationary population of the states, and the arrow width is scaled according to jump probabilities. Additionally, ten randomly selected conformations from each state are overlapped and displayed adjacent to the corresponding node.
@@ III.3.2 HP35
The other 8 states in the 10 state model correspond to states with various degrees of unfolding or misfolded states. Unfolded state 1 in the 5 state model is decomposed into states S1, S2, S3, S4, and S7 in the 10 state model. All these unfolded states contain some degree of secondary structure, but differ in the fraction and location of the structure. Specifically, state S1 presents with alpha helix 1 ($\alpha_{1}$) folded, $\alpha_{2}$ unfolded, and $\alpha_{3}$ misfolded, S2 with $\alpha_{1}$ unfolded and the other two helices misfolded, S3 and S4 with $\alpha_{1}$ unfolded and the other two helices folded, and S7 presenting as a completely unfolded state.
@@
Generally, S7 appears to be a sink for both the misfolded state S1 and the other unfolded states. However, overall, the qualitative analysis of the 10 state model indicates that there is not a dominant folding-unfolding pathway for the HP35 protein; it instead indicates that, even for this simple protein, the folding processes are multifaceted without a single dominant route.
A comparison of the SPIB metastable states at both the 5 and 10 state level of resolution and those discovered by the PCCA+, MPP, and VAMPnet approaches are given in the SI. There we observe that only SPIB and tICA-MPP are able to distinguish clearly and completely the two folded states discriminated by the $\phi_{3}$ dihedral while all other methodologies tested lump those two states into one metastable, folded state at the coarser 5 state level. When the clustering is performed at the finer 10 state level, all methods are able to distinguish the two folded states differing in $\phi_{3}$ angle, although the VAMPnet approach only resolves the two folded states for 40% of the models built. This result is consistent with VAMPnet’s occasional failure in predicting the third-slowest timescale in an alanine dipeptide system[10].